# Quantum Thermal Analogs of Electric Circuits: A Universal Approach

### v3

@@ IV Thermal potentials
where qubits $l$ and $k$ interact with each other via Heisenberg $XX$ type interaction with the interaction strength $J_{lk}$ and $\omega_{k}$ being the transition frequency of the $k$-th qubit. Under the Born-Markov and secular approximations, Eq. (11) with $\widetilde{H}_{S}$ in place of $H_{S}$ dictates the dynamics of the three-qubit system. By equating this to zero, we find the steady-state $\tilde{\rho}^{SS}$ of the three-qubit system. The effective temperatures of the qubits 1, 2, and 3 ($T_{1}$, $T_{2}$, and $T_{3}$, respectively), using the steady-state $\tilde{\rho}^{SS}$ of the three-qubit system are given by
@@
In terms of the elements of the steady state $\tilde{\rho}^{SS}$ of the three-qubit circuit, the quantum thermal currents between the qubits 1 and 3 and between qubits 1 and 2, taking qubit 1 as a node are given by
@@ V Quantum Thermal Diode
Here, we apply the framework developed to a few other quantum thermal circuits, for example, the quantum thermal diode [16, 46] followed by the quantum thermal transistor. The Hamiltonian of the quantum thermal diode is similar to the quantum thermal resistor discussed above, Eq. (10). The difference is that here, we take the temperature of bath $A$ to be non-zero. Accordingly, the dynamical master equation, Eq. (11), changes to accommodate the factor $N_{th,A}=\frac{1}{e^{\omega_{1}/T_{A}}-1}$ with the Lindblad operator $\sigma^{+}_{1}$. The steady-state for this system can be found similarly. We calculate the thermal heat current that flows between bath $A$ and the qubit 1. Based on the temperature difference between bath $A$ and bath $B$, we call the thermal diode to be in the reverse and in the forward bias. The voltage and, correspondingly, the thermal heat currents are zero when the temperatures of both $A$ and $B$ are the same.
Figure 10: Variation of heat current $\mathcal{J}_{A1}$ from bath $A$ to qubit 1 as a function of temperature $T_{A}$ of bath $A$ for a quantum thermal diode. The parameters are: $\omega_{1}=1.5,\omega_{2}=1.5,J_{12}=0.5,\gamma_{A}=0.01,\gamma_{B}=0.1,T_{B}=1.0$.
In Fig. 10, we show the variation of quantum thermal heat current $\mathcal{J}_{A1}$ with temperature of bath $A$. The forward and the reverse bias regions are specified explicitly in the figure. In the forward bias region, the heat current flows from bath $A$ to the qubit 1, whereas in the reverse bias region, the heat current flows from the qubit 1 to the bath $A$. Further, the heat current is zero when the temperatures $T_{A}$ and $T_{B}$ of the baths $A$ and $B$, respectively, are equal.
@@ VI Quantum Thermal Transistor
$+\left.\gamma_{k}N_{th,k}\left[\sigma^{+}_{k}\rho\sigma^{-}_{k}-\frac{1}{2}\left\{\sigma^{-}_{k}\sigma^{+}_{k}\right\}\right]\right),$ (47)
@@ VII.1 The standard quantum thermal Wheatstone bridge
The classical Wheatstone bridge is a device primarily used to determine an unknown resistance. In a circuit of four resistances, one is a tunable resistance, and the other is the unknown, whose value is determined when the ratio between them becomes equal to that of the known resistances. In the case of the quantum Wheatstone bridge, the proposed device would determine the unknown interaction strength of a Hamiltonian. A pictorial comparison between a classical Wheatstone bridge and a quantum one is depicted in [45], wherein a model of quantum Wheatstone bridge was proposed, with the balance condition determining the unknown coupling strength obtained by observing a drop in population of Bell states caused by controlling a tunable coupling. On the other hand, our approach to finding the balance condition is by manipulating the heat current through the circuit, which is in tune with the actual classical Wheatstone bridge.
@@ VIII Details on the results of the quantum thermal adder circuit
Now, using the discussion below Eq. (6) in the manuscript, we can write a relation between the effective temperatures $T_{\alpha},T_{1},T_{2},\dots,T_{n}$ of the qubits $\alpha,1,2,\dots,n$ connected to baths $A,I,II,\dots,N$ (with temperatures $T_{A},T_{I},T_{II},\dots,T_{N}$) and the currents $\mathcal{J}_{A\alpha},\mathcal{J}_{I1},\mathcal{J}_{II2},\dots,\mathcal{J}_{Nn}$ [between (bath, qubit) $=(A,\alpha),(I,1),(II,2),\dots,(N,n)$, respectively] as
@@
for $j=A,I,II,...N$ and $k=\alpha,1,2,...n$. Upon substituting Eq. (51) in (50), keeping in mind $T_{A}=0$, we get

### v4

@@ IV Thermal potentials
where qubits $l$ and $k$ interact with each other via Heisenberg $XX$ type interaction with the interaction strength $J_{lk}$ and $\omega_{k}$ being the transition frequency of the $k$-th qubit. Under the Born-Markov and secular approximations, Eq. (11) with $\widetilde{H}_{S}$ in place of $H_{S}$ dictates the dynamics of the three-qubit system. By equating this to zero, we find the steady-state $\tilde{\rho}^{SS}$ of the three-qubit system. The effective temperatures of the qubits 1, 2, and 3 ($T_{1}$, $T_{2}$, and $T_{3}$, respectively), using the steady-state $\tilde{\rho}^{SS}$ of the three-qubit system, are given by
@@
In terms of the elements of the steady state $\tilde{\rho}^{SS}$ of the three-qubit circuit, the quantum thermal currents between qubits 1 and 3 and between qubits 1 and 2, taking qubit 1 as a node, are given by
@@ V Quantum Thermal Diode
Here, we apply the framework developed to a few other quantum thermal circuits, for example, the quantum thermal diode [16, 46], followed by the quantum thermal transistor. The Hamiltonian of the quantum thermal diode is similar to the quantum thermal resistor discussed above, Eq. (10). The difference is that here, we take the temperature of bath $A$ to be non-zero. Accordingly, the dynamical master equation, Eq. (11), changes to accommodate the factor $N_{th,A}=\frac{1}{e^{\omega_{1}/T_{A}}-1}$ with the Lindblad operator $\sigma^{+}_{1}$. The steady-state for this system can be found similarly. We calculate the thermal heat current that flows between bath $A$ and qubit 1. Based on the temperature difference between baths $A$ and $B$, we determine the reverse and forward bias of the thermal diode. The voltage and, correspondingly, the thermal heat currents are zero when the temperatures of $A$ and $B$ are the same.
Figure 10: Variation of heat current $\mathcal{J}_{A1}$ from bath $A$ to qubit 1 as a function of temperature $T_{A}$ of bath AA for a quantum thermal diode. The parameters are: $\omega_{1}=1.5,\omega_{2}=1.5,J_{12}=0.5,\gamma_{A}=0.01,\gamma_{B}=0.1,T_{B}=1.0$.
In Fig. 10, we show the variation of quantum thermal heat current $\mathcal{J}_{A1}$ with the temperature of bath $A$. The forward and the reverse bias regions are specified explicitly in the figure. In the forward bias region, the heat current flows from bath $A$ to the qubit 1, whereas in the reverse bias region, the heat current flows from the qubit 1 to the bath $A$. Further, the heat current is zero when the temperatures $T_{A}$ and $T_{B}$ of the baths $A$ and $B$, respectively, are equal.
@@ VI Quantum Thermal Transistor
[…]
@@ VII.1 The standard quantum thermal Wheatstone bridge
The classical Wheatstone bridge is a device primarily used to determine an unknown resistance. In a circuit of four resistances, one is a tunable resistance, and the other is the unknown, whose value is determined when the ratio between them becomes equal to that of the known resistances. In the case of the quantum Wheatstone bridge, the proposed device would determine the unknown interaction strength of a Hamiltonian. A pictorial comparison between a classical Wheatstone bridge and a quantum one is depicted in [45], wherein a model of a quantum Wheatstone bridge was proposed, with the balance condition determining the unknown coupling strength obtained by observing a drop in population of Bell states caused by controlling a tunable coupling. On the other hand, our approach to finding the balance condition is by manipulating the heat current through the circuit, which is in tune with the actual classical Wheatstone bridge.
@@ VIII Details on the results of the quantum thermal adder circuit
Now, using the discussion below Eq. (6) in the manuscript, we can write a relation between the effective temperatures $T_{\alpha},T_{1},T_{2},\dots,T_{n}$ of the qubits $\alpha,1,2,\dots,n$ connected to baths $A,I,II,\dots,N$ (with temperatures $T_{A},T_{I},T_{II},\dots,T_{N}$) and the currents $\mathcal{J}_{A\alpha},\mathcal{J}_{I1},\mathcal{J}_{II2},\dots,\mathcal{J}_{Nn}$ [between (bath, qubit) $(A,\alpha),(I,1),(II,2),\dots,(N,n)$, respectively] as
@@
for j=A,I,II,Nj=A,I,II,...N and $k=\alpha,1,2,...n$. Upon substituting Eq. (51) in (50), keeping in mind $T_{A}=0$, we get